# A Framework for Lorentz-Dirac Dynamics and Post-Newtonian Interaction of Radiating Point Charges

### v4

@@ II.2.2 Mechanical Momentum and Mass Renormalization
We now use this identity to rewrite the self-force term in Eq. (16). Substituting into Eq. (16) and projecting orthogonally to the four-velocity gives the standard covariant Lorentz–Dirac equation,
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta}.$ (20)
Equation (20) is third order in the particle trajectory, since it contains $\dot{a}^{\alpha}=da^{\alpha}/d\tau$. This leads to runaway and preaccelerated solutions. The physically meaningful equation governing radiation reaction is obtained by applying the reduction-of-order prescription of Landau and Lifshitz [10, 13], replacing $\dot{a}^{\alpha}$ by the proper-time variation of the external field.
III Resolution via Reduction of Order
The Lorentz–Dirac equation,
$ma^{\alpha}=F^{\alpha}_{\mathrm{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)\dot{a}^{\beta},$ (21)
is third order in the worldline $z^{\alpha}(\tau)$, since it involves the derivative of the acceleration $\dot{a}^{\alpha}$. The requirement to specify initial position, velocity, and acceleration leads to well-known difficulties: runaway solutions and preacceleration. These features are already clearly visible in the nonrelativistic limit [8]. In the nonrelativistic regime, Eq. (21) reduces to
$a(t)-t_{0}\,\dot{a}(t)=\frac{1}{m}F_{\mathrm{ext}}(t),\qquad t_{0}\equiv\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}m}\$(22)
Considering a force that is switched on abruptly and held constant,
$F_{\mathrm{ext}}(t)=f\,\theta(t),$ (23)
where $\theta(t)$ is the Heaviside step function. Solving Eq. (22) gives
$a(t)=e^{t/t_{0}}\!\left[b-\frac{f}{m}\bigl(1-e^{-t/t_{0}}\bigr)\theta(t)\right].$ (24)
For a generic constant $b$, the exponential factor dominates for $t\gg t_{0}$,
$a(t)\sim b\,e^{t/t_{0}},$ (25)
so that even in the presence of a constant applied force the acceleration grows without bound. These unphysical runaway solutions occur whenever the third-derivative term is retained. They can be eliminated only by choosing the constant
$b=\frac{f}{m},$
which removes the runaway term in (24). With this choice, we find
$a(t)=\frac{f}{m}\!\left[\theta(-t)\,e^{t/t_{0}}+\theta(t)\right].$ (26)
Equation (26) shows that the particle begins to accelerate before the force is applied. Although $F_{\mathrm{ext}}(t)=0$ for $t<0$, the acceleration is nonzero for $t<0$:
$a(t)=\frac{f}{m}e^{t/t_{0}},\qquad t<0.$
This phenomenon is known as preacceleration. The particle “anticipates” the force a time $t_{0}$ in advance. For an electron,
$t_{0}=\frac{2}{3}\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}m}\approx 6\times 10^{-24}\,\mathrm{s},$
so preacceleration occurs only on extremely small timescales; nevertheless, its appearance indicates a conceptual inconsistency in treating the charge as a strict point particle.
These difficulties are resolved by treating the radiation-reaction term as a small correction and performing reduction of order [10, 13]. This yields the modified Dirac-Lorentz equations, which is second order, causal, and free of runaway behaviour. Equation (20) is third order in time and possesses unphysical runaway and preaccelerated solutions. Treating $q^{2}/m$ as small and replacing $\dot{a}^{\alpha}$ using the leading-order equation of motion gives
$ma^{\alpha}=F^{\alpha}_{\text{ext}}+\frac{2}{3}\,\frac{q^{2}}{4\pi\varepsilon_{0}c^{3}m}\,\left(\delta^{\alpha}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}}\right)F^{\beta}_{\text{ext},\gamma}u^{\gamma},$ (27)
the Landau–Lifshitz form, which is second-order and free of pathologies. We first consider
$ma^{\alpha}$ $=F^{\alpha}_{\rm ext}+t_{0}\,(\delta^{\alpha}_{\ \beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}})\,F_{\rm ext}^{\beta}{}_{,\gamma}\,u^{\gamma},$ (28)
where $u^{\alpha}$ is the four-velocity, $a^{\alpha}=du^{\alpha}/d\tau$ the four-acceleration, and $t_{0}$ the small radiation-reaction timescale. For an electromagnetic external field we have
$F^{\alpha}_{\rm ext}=q\,F_{\rm ext}^{\alpha\mu}\,u_{\mu},$ (29)
with $F_{\rm ext}^{\alpha\mu}$ the field tensor. Computing the derivative of $F_{\rm ext}^{\beta}$ along the worldline:
$F_{\rm ext}^{\beta}{}_{,\gamma}u^{\gamma}$ $=q\left[(\partial_{\gamma}F_{\rm ext}^{\beta\mu})\,u_{\mu}u^{\gamma}+F_{\rm ext}^{\beta\mu}\,(\partial_{\gamma}u_{\mu})\,u^{\gamma}\right].$ (30)
Since $(\partial_{\gamma}u_{\mu})\,u^{\gamma}=du_{\mu}/d\tau=a_{\mu}$, this becomes
$F_{\rm ext}^{\beta}{}_{,\gamma}u^{\gamma}$ $=q\,F_{\rm ext}^{\beta\mu}{}_{,\gamma}\,u_{\mu}u^{\gamma}+q\,F_{\rm ext}^{\beta\mu}\,a_{\mu}.$ (31)
$ma^{\alpha}$ $=q\,F_{\rm ext}^{\alpha\mu}\,u_{\mu}+qt_{0}\,(\delta^{\alpha}{}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}})\,F_{\rm ext}^{\beta\mu}{}_{,\gamma}\,u_{\mu}u^{\gamma}$
$\quad+qt_{0}\,(\delta^{\alpha}{}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}})\,F_{\rm ext}^{\beta\mu}\,a_{\mu}.$ (32)
Inside the $O(t_{0})$ term we replace $a_{\mu}$ by its leading-order value from the Lorentz force:
$a_{\mu}$ $\approx\frac{q}{m}\,F_{{\rm ext}\,\mu\nu}\,u^{\nu}.$ (33)
Thus,
$F_{\rm ext}^{\beta\mu}\,a_{\mu}$ $\approx\frac{q}{m}\,F_{\rm ext}^{\beta\mu}\,F_{{\rm ext}\,\mu\nu}\,u^{\nu}.$ (34)
Putting everything together (one term is zero by antisymmetry):
$ma^{\alpha}$ $=q\,F_{\rm ext}^{\alpha}{}_{\mu}\,u^{\mu}+qt_{0}\,F_{\rm ext}^{\alpha}{}_{\mu,\nu}\,u^{\mu}u^{\nu}$
$\quad+\frac{q^{2}t_{0}}{m}\,(\delta^{\alpha}{}_{\beta}+\frac{u^{\alpha}u_{\beta}}{c^{2}})\,F_{\rm ext}^{\beta}{}_{\mu}\,F_{\rm ext}^{\mu}{}_{\nu}\,u^{\nu}.$ (35)
This is the desired reduced-order Landau–Lifshitz form of the Lorentz–Dirac equation and will be a central equation for our single particle simulations of radiation reaction. Expanding it out in full
$ma^{\alpha}=q\,F_{\rm ext}^{\alpha}{}_{\mu}\,u^{\mu}+q\,t_{0}\!\left[F_{\rm ext}^{\alpha}{}_{\mu,\nu}\,u^{\mu}u^{\nu}+\frac{q}{m}\,F_{\rm ext}^{\alpha}{}_{\mu}\,F_{\rm ext}^{\mu}{}_{\nu}\,u^{\nu}+\frac{q}{mc^{2}}\,u^{\alpha}\!\left(u_{\beta}\,F_{\rm ext}^{\beta}{}_{\mu}\,F_{\rm ext}^{\mu}{}_{\nu}\,u^{\nu}\right)\right],$ (36)

### v5

@@ II Radiation reaction: Lorentz–Dirac, Landau–Lifshitz reduction, and the near-zone limit
In the near zone ($v\ll c$), dissipative electromagnetic radiation reaction first enters at 1.51.5PN order and is controlled by the electric dipole moment
$\mathbf{d}(t)=\sum_{a}q_{a}\,\mathbf{x}_{a}(t).$ (10)
To leading order, the 1.51.5PN near–zone field gives the dipole radiation–reaction force
$\mathbf{F}_{a}^{(1.5{\rm PN})}=\frac{q_{a}}{4\pi\varepsilon_{0}}\frac{2}{3c^{3}}\,\dddot{\mathbf{d}}(t),$ (11)
and the associated mechanical energy loss equals the dipole (Larmor) power,
$\frac{dE}{dt}=-\frac{1}{4\pi\varepsilon_{0}}\frac{2}{3c^{3}}\,|\ddot{\mathbf{d}}(t)|^{2},$ (12)
up to a total time derivative (Schott energy). In our single–particle tests we verify the covariant energy balance relation
$\Delta K=W_{\rm ext}-E_{\rm rad},$ (13)
and in the remainder of this paper we adopt the order–reduced LL dynamics as the dissipative input for the post–Newtonian many–body framework developed below.
III The N-body Darwin–radiation-reaction system
We construct the electromagnetic NN–body dynamics in direct analogy with the PN/EOB treatment of compact binaries: a conservative Hamiltonian sector through 11PN ($\mathcal{O}(c^{-2})$), generated by the Darwin Hamiltonian, augmented by a non-Hamiltonian 1.51.5PN dipole radiation–reaction force ($\mathcal{O}(c^{-3})$). We work in SI units, with $k=(4\pi\varepsilon_{0})^{-1}$ and $\mathbf{r}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b},\;r_{ab}=|\mathbf{r}_{ab}|,\;\hat{\mathbf{r}}_{ab}=\mathbf{r}_{ab}/r_{ab}.$
The conservative dynamics is generated by the Darwin Hamiltonian
$\quad-\frac{1}{2}\sum_{a\neq b}k\,\frac{q_{a}q_{b}}{2m_{a}m_{b}c^{2}}\,\frac{1}{r_{ab}}\!\left[\mathbf{p}_{a}\!\cdot\!\mathbf{p}_{b}+\frac{(\mathbf{p}_{a}\!\cdot\!\mathbf{r}_{ab})(\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab})}{r_{ab}^{2}}\right].$ (14)
The leading dissipative contribution is the dipole radiation–reaction force, evaluated by order reduction and expressed entirely in canonical variables,
$\mathbf{F}_{a}^{(1.5\mathrm{PN})}(t)=\frac{1}{(4\pi\varepsilon_{0})^{2}}\frac{2q_{a}}{3c^{3}}\sum_{b}\sum_{c\neq b}\frac{q_{b}^{2}q_{c}}{m_{b}r_{bc}^{3}}\Big[\tfrac{\mathbf{p}_{b}}{m_{b}}-\tfrac{\mathbf{p}_{c}}{m_{c}}$
$-3\hat{\mathbf{r}}_{bc}\big(\hat{\mathbf{r}}_{bc}\!\cdot\!(\tfrac{\mathbf{p}_{b}}{m_{b}}-\tfrac{\mathbf{p}_{c}}{m_{c}})\big)\Big],$ (15)
which is fully consistent with electromagnetic energy balance.
Collecting results, the closed 11PN+1.5+1.5PN equations of motion in canonical phase space are
$\dot{\mathbf{x}}_{a}$ $=\frac{\partial H_{\rm D}}{\partial\mathbf{p}_{a}}=\frac{\mathbf{p}_{a}}{m_{a}}-\frac{\mathbf{p}_{a}^{2}\,\mathbf{p}_{a}}{2m_{a}^{3}c^{2}}$
$\quad-\sum_{b\neq a}\frac{q_{a}q_{b}}{8\pi\varepsilon_{0}}\frac{1}{m_{a}m_{b}c^{2}}\frac{1}{r_{ab}}\left[\mathbf{p}_{b}+\mathbf{r}_{ab}\frac{\mathbf{p}_{b}\!\cdot\!\mathbf{r}_{ab}}{r_{ab}^{2}}\right],$
with the conservative force written explicitly as
$-\frac{\partial H_{\rm D}}{\partial\mathbf{x}_{a}}$ $=\sum_{b\neq a}k\,\frac{q_{a}q_{b}}{r_{ab}^{2}}\,\hat{\mathbf{r}}_{ab}$ (17)
$\quad+k\sum_{b\neq a}\frac{q_{a}q_{b}}{2c^{2}r_{ab}^{2}}\Bigg[\frac{\mathbf{p}_{a}}{m_{a}}\left(\hat{\mathbf{r}}_{ab}\!\cdot\!\frac{\mathbf{p}_{b}}{m_{b}}\right)+\frac{\mathbf{p}_{b}}{m_{b}}\left(\hat{\mathbf{r}}_{ab}\!\cdot\!\frac{\mathbf{p}_{a}}{m_{a}}\right)$
$\qquad\qquad-\hat{\mathbf{r}}_{ab}\left(\frac{\mathbf{p}_{a}\!\cdot\!\mathbf{p}_{b}}{m_{a}m_{b}}+3\left(\hat{\mathbf{r}}_{ab}\!\cdot\!\frac{\mathbf{p}_{a}}{m_{a}}\right)\left(\hat{\mathbf{r}}_{ab}\!\cdot\!\frac{\mathbf{p}_{b}}{m_{b}}\right)\right)\Bigg].$
Equations (16)–(17), together with (15), constitute the fully explicit and implementable 11PN+1.5+1.5PN phase–space system used in our simulations. When $\mathbf{F}_{a}^{(1.5\mathrm{PN})}$ is switched off, the evolution is strictly conservative and exhibits no secular drift in $H_{\rm D}$, providing a stringent validation of the Darwin sector prior to the inclusion of dissipative radiation reaction.
IV Specializing to charged binaries: $N=2$
We now present a brief analysis of the electromagnetic radiation–reaction term used Eq.(15) in our $N$-body framework, specialized to $N=2$ and connect it explicitly to both the Einstein–Maxwell post–Newtonian system of Placidi et al. [36] and the analysis of Kunze and Spohn [35]. We start from the many-body electromagnetic radiation–reaction force at 1.51.5PN order, written in canonical variables, Eq. (15) and specialize to a binary system with particle labels $a,b,c\in\{1,2\}$ and focus on the radiation–reaction force acting on particle $1$. The double sum contains exactly two nonvanishing contributions, corresponding to the ordered pairs $(b,c)=(1,2)$ and $(b,c)=(2,1)$.
For $(b,c)=(1,2)$ we obtain
$\mathbf{a}^{\mathrm{RR}}_{1}=\frac{2q_{1}^{2}q_{2}}{3(4\pi\varepsilon_{0})^{2}c^{3}}\frac{1}{m_{1}r^{3}}\left(\frac{q_{1}}{m_{1}}-\frac{q_{2}}{m_{2}}\right)\left[\mathbf{v}-3\hat{\mathbf{r}}\big(\hat{\mathbf{r}}\!\cdot\!\mathbf{v}\big)\right].$ (20)
where $\hat{\mathbf{r}}\equiv\hat{\mathbf{r}}_{12}$. Equation (20) exhibits the characteristic electromagnetic dipole radiation–reaction structure and is proportional to the difference of charge-to-mass ratios. Consequently, when $q_{1}/m_{1}=q_{2}/m_{2}$ the electromagnetic dipole moment is constant and the 1.51.5PN radiation–reaction force vanishes identically. This behavior is observed in the numerical simulations and in agreement with the result of Kunze and Spohn, [35]. On a related note, Placidi et al.  [36] recently derived the 1.51.5PN dissipative acceleration for charged binaries in Einstein–Maxwell theory using Hadamard regularization. Taking the limit $G\to 0$ which isolates the purely electromagnetic contributions, we obtain perfect agreement with their results. The more general formula in Eq. (15) holds for an N-body system of point charges. At leading dissipative order (1.5PN), the radiation-reaction acceleration for the two-body electromagnetic system reduces in relative coordinates to the compact form
[…]